Use data in the following table to explain the economic effects of a price ceiling at $6, at $5, and at $4.
Quantity Quantity
Price demanded supplied
$7 4,500 4,500
6 5,000 3,500
5 5,500 2,500
4 6,000 1,500


Use data in the following table to explain the economic effects of a price floor at $8, at $9, and at $10. Explain the economic effects.

Quantity Quantity
Price demanded supplied
$10 3,000 7,500
9 3,500 6,500
8 4,000 5,500
7 4,500 4,500
1
Expert's answer
2020-06-03T10:34:02-0400

Question1: Use data in the following table to explain the economic effects of a price ceiling at $6, at $5, and at $4.


Price Demand Supply

$7 4,500 4,500

$6 5,000 3,500

$5 5,500 2,500

$4 6,000 1,500


For a price ceiling to be effective, it must be set below the equilibrium price.


From the data, the equilibrium price is $7 and the equilibrium quantity that is demanded and supplied in the market is 4,500.


If a price ceiling is set at $6, there will be a shortage in the market since the quantity demanded is greater than the quantity supplied in the market. A shortage of 5,000 - 3,500 = 1,500 will be created.


If the price ceiling is set at $5, a shortage of 5,500 - 2,500 = 3,000 will be created.


If the price ceiling is set at $4, a shortage of 5,000 - 1,500 = 3,500 will be created.


Question 2: Use data in the following table to explain the economic effects of a price floor at $8, at $9, and at $10. Explain the economic effects. 



Price Demand Supply

$10 3,000 7,500

$9 3,500 6,500

$8 4,000 5,500

$7 4,500 4,500


From the information, the equilibrium price is $7 and the equilibrium quantity demanded and supplied in the market is 4,500.


For a price floor to be effective, it must be set above the equilibrium price.


At a price floor of $8, the quantity supplied will be greater than the quantity demanded. A surplus of 5,500 - 4,000 = 1,500 will be created.


At a price floor of $9, a surplus of 6,000 - 3,500 = 3,000 will be created.


At a price floor of $10, a surplus of 7,500 - 3,000 = 4,000 will be created.